### v1

# Large Language Model Compression with Global Rank and Sparsity Optimization

Abstract
Low-rank and sparse composite approximation is a natural idea to compress Large Language Models (LLMs). However, such an idea faces two primary challenges that adversely affect the performance of existing methods. The first challenge relates to the interaction and cooperation between low-rank and sparse matrices, while the second involves determining weight allocation across different layers, as redundancy varies considerably among them. To address these challenges, we propose a novel two-stage LLM compression method with the capability of global rank and sparsity optimization. It is noteworthy that the overall optimization space is vast, making comprehensive optimization computationally prohibitive. Therefore, to reduce the optimization space, our first stage utilizes robust principal component analysis to decompose the weight matrices of LLMs into low-rank and sparse components, which span the low dimensional and sparse spaces containing the resultant low-rank and sparse matrices, respectively. In the second stage, we propose a probabilistic global optimization technique to jointly identify the low-rank and sparse structures within the above two spaces. The appealing feature of our approach is its ability to automatically detect the redundancy across different layers and to manage the interaction between the sparse and low-rank components. Extensive experimental results indicate that our method significantly surpasses state-of-the-art techniques for sparsification and composite approximation.
1 Introduction
Transformer-based large language models (LLMs) Vaswani et al. [2023], Touvron et al. [2023b], OpenAI et al. [2024] have achieved remarkable progress across natural language processing (NLP), computer vision, multimodal tasks, and scientific applications. Despite these successes, their massive parameter sizes pose critical challenges: they demand huge storage and memory footprints, incur slow inference speeds, and require substantial computational resources for training. Consequently, model compression Cheng et al. [2020], Wang et al. [2024], Zhu et al. [2024] has become an essential line of research for enabling real-world LLM deployment under stringent hardware constraints.
Among compression strategies, quantization Han et al. [2015], Chee et al. [2023], Kuzmin et al. [2023] typically retains overall model structure by reducing the precision of weights, thus often preserving performance. By contrast, pruning Liu et al. [2017], Frankle and Carbin [2019], Sun et al. [2023], Frantar and Alistarh [2023] removes individual weights based on certain criteria (e.g., magnitude or importance scores). Although pruning is flexible and can yield substantial parameter savings, it may degrade performance unless combined with additional fine-tuning or distillation Sanh et al. [2020], especially in large-scale LLMs that encode extensive linguistic and factual knowledge Geva et al. [2021], Dai et al. [2022].
To retain more critical information under aggressive compression, researchers have explored “low-rank plus sparse" decompositions Li et al. [2023], Ren and Zhu [2023], Han et al. [2024]. In this approach, the weight matrix is decomposed into a low-rank part that captures global correlations and a sparse part that highlights outliers or domain-specific knowledge. However, existing methods often rely on manually set singular-value thresholds, which can inadvertently discard medium-sized yet important singular values. Additionally, these methods frequently require computationally expensive backpropagation for parameter updates. While there is some interaction between the optimization of the low-rank and sparse components, the two parts are still relatively independent in their update processes. Lastly, due to the significant redundancy variations from early layers to deeper ones, how to allocate rank and sparsity across layers in a globally optimal manner remains unclear.
In this paper, we address these issues via a novel two-stage compression framework tailored to LLMs. First, we apply robust principal component analysis (RPCA) Candes et al. [2009] to factor each weight matrix into strictly low-rank and sparse components, thereby reducing the otherwise huge search space into a low-dimensional subspace and a sparse subspace. Second, we introduce a probabilistic global optimization scheme that jointly determines which singular values in the low-rank component and which nonzero entries in the sparse component should be retained. This is done by assigning Bernoulli probabilities and updating them via policy gradient Williams [1992] on a small calibration set, avoiding heuristic thresholds or large-scale gradient updates. Critically, our method automatically detects the differing redundancy levels across layers and manages the interaction between low-rank and sparse parts, ensuring that vital parameters are kept while truly redundant ones are pruned away. We summarize our main contributions as follows:
  • We propose a two-stage LLM compression approach that first uses RPCA to produce low-rank and sparse subspaces, then employs a Bernoulli-based global optimization for rank and sparsity selection.
  • Our framework eliminates the need for manual thresholds or layerwise iterative backpropagation, offering an end-to-end scheme that adapts automatically to various layers’ redundancy characteristics.
  • Extensive experiments show that our method outperforms existing sparsification and composite approximation baselines under multiple compression ratios, highlighting its effectiveness and robustness.
2 Related Work
2.1 Unstructured Pruning
Unstructured pruning eliminates individual weights by setting them to zero, providing fine-grained control over model sparsity. SparseGPT Frantar and Alistarh [2023] leverages second-order information to perform layer-wise pruning with minimal retraining, whereas Wanda Sun et al. [2024] combines weight magnitude with activation statistics for a more straightforward pruning strategy. While these methods achieve efficient pruning, they struggle to maintain performance at high sparsity levels and often require additional retraining Sanh et al. [2020], Renda et al. [2020]. BESA Xu et al. [2024] introduces a differentiable pruning framework that dynamically allocates sparsity across layers to minimize performance degradation, producing competitive results without requiring extensive retraining. Dynamic Sparse Training (DST) Liu et al. [2020] proposes an end-to-end sparse training method where trainable pruning thresholds dynamically adjust the sparsity level during training. Unlike post-training pruning methods, DST eliminates the need for iterative fine-tuning by continuously optimizing layer-wise sparsity using backpropagation. DST is designed for training sparse networks from scratch.
2.2 Low-Rank Plus Sparse
Low-rank matrix approximation is a widely adopted approach for reducing model size. By decomposing weight matrices into reduced-rank components via SVD, low-rank approximation preserves essential information while significantly lowering memory and computational costs Denton et al. [2014], Li et al. [2023]. Recent works such as LoSparse Li et al. [2023] combine low-rank decomposition with sparse corrections to improve compression efficiency. However, these methods typically rely on heuristic thresholds and necessitate iterative fine-tuning. LPAF Ren and Zhu [2023] extends this concept by first performing structured pruning to obtain a low-rank sparse model, followed by a sparsity-aware SVD decomposition of the weight matrices. A mixed-rank fine-tuning process then reconstructs the model, leading to a compressed structure where $W\approx AB$. While effective, these methods still require multiple post-processing steps to maintain performance. Beyond post-hoc compression techniques, some works explore model sparsification during pretraining. SLTrainHan et al. [2024] formulates the sparse-plus-low-rank objective directly in the pretraining stage, jointly learning sparse and low-rank representations—a different paradigm from post-hoc compression. Meanwhile, SLIM Mozaffari and Dehnavi [2024] proposes a one-shot quantized sparse plus low-rank framework that eschews retraining by combining symmetric quantization with saliency-based decomposition. However, SLIM primarily focuses on quantization-aware sparsity. By comparison, our approach directly compresses pre-trained models using robust principal component analysis (RPCA) to globally identify both low-rank and sparse structures, thus avoiding ad-hoc thresholds and extensive re-training.
3 Preliminary
Figure 1: Singular values of Layer 0 and Layer 31 across different modules, comparing original and RPCA-processed matrices. The dotted line represents the singular value distribution of the original model, and the solid line represents the singular value distribution of the low-rank matrix after RPCA processing.
4 Preliminaries
4.1 Low-Rank Approximation
Low-rank approximation Chu et al. [2003] is a fundamental technique in matrix theory, widely used to reduce the parameter count in neural networks while preserving most of the model’s performance. In large language models (LLMs), weight matrices are typically high-dimensional and dense. By approximating a weight matrix $\mathbf{W}\in\mathbb{R}^{M\times N}$ as $\mathbf{U}\mathbf{V}^{\top}$ with rank $R\ll\min(M,N)$, we can achieve substantial reductions in storage and computational cost. Concretely, one typically uses Singular Value Decomposition (SVD) to write
$\mathbf{W}=\mathbf{U}\boldsymbol{\Sigma}\mathbf{V}^{\top},$ (1)
and then retains only the largest $R$ singular values $\boldsymbol{\Sigma}_{R}$, yielding
$\mathbf{W}\,\approx\,\mathbf{U}_{R}\,\boldsymbol{\Sigma}_{R}\,\mathbf{V}_{R}^{\top}.$ (2)
This factorization can reduce the parameter count from $M\times N$ to $(M+N)\times R$, and also break a large matrix multiplication into smaller ones:
$\mathbf{W}\mathbf{x}\;\approx\;\mathbf{U}_{R}\bigl{(}\boldsymbol{\Sigma}_{R}(\mathbf{V}_{R}^{\top}\mathbf{x})\bigr{)},$
leading to efficiency gains.
Despite these benefits, low-rank approximation alone may be insufficient for LLM compression, especially when the singular values do not decay sharply. Figures 1 illustrate the singular value distributions for two layers (Layer 0 and Layer 31) in a large Transformer. The dashed lines represent the original matrices, showing that certain modules in the same Transformer block (e.g., attention vs. feedforward) might exhibit similar shapes, yet across different layers, the redundancy patterns can vary considerably. Consequently, imposing the same rank $R$ uniformly across all layers may prune too aggressively in some places and insufficiently in others. A more flexible approach is needed to handle these differences among modules and layers.
4.2 Low-Rank Approximation with Sparse Corrections
To mitigate the shortcomings of purely low-rank approximation, recent methods Li et al. [2023], Ren and Zhu [2023] advocate combining a low-rank matrix with a sparse correction term. One splits the model weights as:
$\mathbf{W}\;=\;\underbrace{\mathbf{U}\mathbf{V}^{\top}}_{\text{low-rank}}\;+\;\underbrace{\mathbf{S}}_{\text{sparse}}.$ (3)
LoSparse Li et al. [2023], for example, first applies an SVD on $\mathbf{W}$ to obtain a low-rank component (with some rank $R$), and then prunes the residual $\mathbf{W}-\mathbf{U}\mathbf{V}^{\top}$ to form a sparse matrix $\mathbf{S}$. In practice, one must still decide the singular-value cutoff (or target rank) and the sparsity ratio for the residual. Often, additional fine-tuning is performed on the low-rank part to recover lost performance, or iterative pruning is applied to the sparse part Molchanov et al. [2019], which can be computationally expensive.
A major limitation of these approaches is that they rely heavily on manually specified thresholds for both singular values and residual pruning. They also lack a clear mechanism to coordinate how much rank vs. how much sparsity each layer should receive, since different layers and modules may have different redundancy patterns. Furthermore, when both the low-rank matrix and the sparse matrix need simultaneous updates (or fine-tuning), memory consumption can become large, often exceeding the budget for fine-tuning.
Figure 2: Overview of our proposed compression method. The weight matrix $\mathbf{W}$ is decomposed into a low-rank component $\mathbf{L}$ and a sparse component $\mathbf{S}$ using RPCA. Both components are pruned through Bernoulli sampling guided by learned probability scores, optimized via policy gradient. The low-rank component is further factorized into $\mathbf{U}^{\prime}$ and $\mathbf{V}^{\prime}$ to reduce the number of model para meters.
5 Method
We begin by formulating the global objective of compressing LLM weights under a parameter budget (§ 5.1). We then describe our proposed approach (§ 5.2), which first uses RPCA to decompose each weight matrix into low-rank and sparse components, and subsequently prunes these components in a probabilistic manner, without heuristic thresholds or large-scale fine-tuning. Additional theoretical analysis can be found in the Appendix F.
5.1 Problem Formulation
Suppose we have $L$ layers in an LLM, each containing weight matrices $\{\mathbf{W}^{(l)}\}_{l=1}^{L}$. We seek compressed matrices $\{\tilde{\mathbf{W}}^{(l)}\}$ such that the total parameter count does not exceed a budget $K$, while minimizing a loss $\ell(\tilde{\mathbf{W}})$ measured on a small calibration set $\mathcal{D}$. Formally,
$\min_{\{\tilde{\mathbf{W}}^{(l)}\}}\quad\sum_{(x,y)\in\mathcal{D}}\ell\Bigl{(}f(\tilde{\mathbf{W}};x),\,y\Bigr{)},$ (4)
subject $\text{subject to}\quad\mathrm{ParamCount}\bigl{(}\{\tilde{\mathbf{W}}^{(l)}\}\bigr{)}\;\leq\;K,$
where $f(\tilde{\mathbf{W}};x)$ is the LLM’s forward pass given the compressed weights, and $\mathrm{ParamCount}(\cdot)$ measures how many parameters are retained. Directly pruning each individual weight is intractable for very large matrices. To address this, we propose to:
  • Decompose each $\mathbf{W}^{(l)}$ via RPCA to obtain a low-rank matrix $\mathbf{L}$ and a sparse matrix $\mathbf{S}$, reducing the search space to “global rank directions” plus “sparse outliers.”
  • Probabilistically prune both components under the budget $K$ by learning Bernoulli retention probabilities through policy gradient on a small calibration set.
5.2 Proposed Approach: CAP
As illustrated in Figure 2, we present an overview of our proposed method. Each weight matrix $\mathbf{W}$ in the model is decomposed into two components: a low-rank component $\mathbf{L}$, which captures dominant patterns, and a sparse component $\mathbf{S}$, which represents more fine-grained information. This representation enables the compression of the model while preserving its performance. To achieve this decomposition, we leverage RPCA. After decomposition, we selectively prune $\mathbf{L}$ and $\mathbf{S}$ by modeling their parameters using Bernoulli distributions, which are optimized via policy gradient. This process achieves adaptive pruning without relying on heuristic thresholds or global fine-tuning. In the following sections, we provide detailed explanations of our algorithm.
5.2.1 Robust Subspace Decomposition
The first stage aims to construct separated subspaces for low-rank and sparse components, addressing two critical challenges: (1) avoiding heuristic rank selection through automated subspace decomposition, and (2) establishing proper interaction mechanisms between structural patterns and sparse residuals. We achieve this through RPCA, which decomposes each weight matrix $\mathbf{W}\in\mathbb{R}^{m\times n}$ as:
$\mathbf{W}\;=\;\mathbf{L}+\mathbf{S}$ (5)
This decomposition fundamentally differs from conventional low-rank approximations by simultaneously addressing three key requirements: automatic rank determination, avoiding preset rank thresholds through nuclear norm optimization; outlier preservation, maintaining performance-critical sparse elements via $\ell_{1}$-constrained sparsity; and subspace orthogonality, ensuring the low-rank ($\mathbf{L}$) and sparse ($\mathbf{S}$) components inhabit complementary subspaces.
The unified RPCA optimization framework achieves these objectives through:
$\min_{\mathbf{L},\mathbf{S}}\underbrace{\|\mathbf{L}\|_{*}}_{\text{Low-rank constraint}}+\lambda\,\underbrace{\|\mathbf{S}\|_{1}}_{\text{Sparsity constraint}}\text{subject to}\mathbf{W}=\mathbf{L}+\mathbf{S},$ (6)
Where the nuclear norm $\|\mathbf{L}\|_{*}$ (sum of singular values) automatically determines the effective rank through singular value shrinkage, while the $\ell_{1}$ penalty on $\mathbf{S}$ provides adaptive outlier selection by suppressing small perturbations while preserving significant deviations. The convex optimization structure guarantees global convergence, ensuring stable decomposition across different network layers.
We implement this through an Augmented Lagrangian Method (ALM) with Alternating Direction Method of Multipliers (ADMM) Lin et al. [2010], which iteratively updates:
$\mathbf{L}_{k+1}$ $=\arg\min_{\mathbf{L}}\ \|\mathbf{L}\|_{*}$
$\quad+\frac{\mu}{2}\left\|\mathbf{W}-\mathbf{L}-\mathbf{S}_{k}+\mu^{-1}\mathbf{Y}_{k}\right\|_{F}^{2},$ (7)
$\mathbf{S}_{k+1}$ $=\arg\min_{\mathbf{S}}\ \lambda\|\mathbf{S}\|_{1}$
$\quad+\frac{\mu}{2}\left\|\mathbf{W}-\mathbf{L}_{k+1}-\mathbf{S}+\mu^{-1}\mathbf{Y}_{k}\right\|_{F}^{2},$ (8)
$\mathbf{Y}_{k+1}$ $=\mathbf{Y}_{k}+\mu\left(\mathbf{W}-\mathbf{L}_{k+1}-\mathbf{S}_{k+1}\right).$ (9)
The $\mathbf{L}$-update employs Singular Value Thresholding (SVT) Cai et al. [2008]:
$\mathbf{L}_{k+1}=\mathbf{U}\operatorname{diag}(\operatorname{shrink}_{\mu^{-1}}(\boldsymbol{\sigma}))\mathbf{V}^{\top}$ (10)
where $\mathbf{U}\boldsymbol{\sigma}\mathbf{V}^{\top}$ is the SVD of $\mathbf{W}-\mathbf{S}_{k}+\mu^{-1}\mathbf{Y}_{k}$, with singular value shrinkage $\operatorname{shrink}_{\tau}(\sigma_{i})=\max(\sigma_{i}-\tau,0)$. The $\mathbf{S}$-update applies elementwise soft-thresholding:
$[\mathbf{S}_{k+1}]_{ij}=\operatorname{shrink}_{\lambda\mu^{-1}}([\mathbf{W}-\mathbf{L}_{k+1}+\mu^{-1}\mathbf{Y}_{k}]_{ij})$ (11)
This alternating optimization progressively separates the weight matrix into a low-dimensional subspace capturing directional patterns ($\mathbf{L}$) and a sparse subspace containing localized refinements ($\mathbf{S}$), establishing the foundation for subsequent global resource allocation.
5.2.2 Learnable Probabilistic Pruning
In the second stage, we jointly refine important parameters in both spaces via a probabilistic global optimization technique. It is able to automatically detect redundancies between different layers and manage interactions between sparse and low-rank components without requiring computationally expensive back-propagation.
Having decomposed $\mathbf{W}$ into $\mathbf{L}$ (with singular values $\boldsymbol{\sigma}=(\sigma_{1},\dots,\sigma_{r})$) and a sparse part $\mathbf{S}$, we next decide which rank-$1$ directions in $\mathbf{L}$ and which entries in $\mathbf{S}$ to keep.
Each retained singular value $\sigma_{i}$ requires storing its corresponding singular vectors $\mathbf{u}_{i}\in\mathbb{R}^{m}$ and $\mathbf{v}_{i}\in\mathbb{R}^{n}$, contributing $(m+n)$ parameters. Denote the nonzero entries of $\mathbf{S}$ by $\{s_{ij}\}$. We introduce Bernoulli random variables:
$m_{\sigma_{i}}\;\sim\;\mathrm{Bernoulli}(s_{\sigma_{i}}),\quad m_{S_{ij}}\;\sim\;\mathrm{Bernoulli}(s_{S_{ij}}),$
where $s_{\sigma_{i}}\in[0,1]$ and $s_{S_{ij}}\in[0,1]$ are learned retention probabilities. The compressed matrix is then
$\tilde{\mathbf{W}}\;=\;\mathbf{U}\,\mathrm{diag}\!\bigl{(}\boldsymbol{\sigma}\odot\mathbf{m}_{\sigma}\bigr{)}\,\mathbf{V}^{\top}\;+\;\mathbf{S}\,\odot\,\mathbf{m}_{S},$ (12)
subject to $\,\sum_{i}s_{\sigma_{i}}(m+n)\;+\;\sum_{i,j}s_{S_{ij}}\,\leq\,K\,$ to respect the total parameter budget.
Learning probabilities via policy gradient.
We minimize the expected loss on a small calibration set $\mathcal{D}$:
$\min_{\mathbf{s}}\ \mathbb{E}_{\mathbf{m}\sim p(\mathbf{m}\mid\mathbf{s})}\Bigl{[}\,\mathcal{L}(\tilde{\mathbf{W}})\,\Bigr{]},$ (13)
where $\mathbf{s}=\{s_{\sigma_{i}},s_{S_{ij}}\}$ and $p(\mathbf{m}\mid\mathbf{s})$ is the product of Bernoulli distributions. We employ a REINFORCE-style Williams [1992] policy gradient:
$\nabla_{s_{k}}\,\mathbb{E}_{\mathbf{m}}[\mathcal{L}(\tilde{\mathbf{W}})]\;=\;\mathbb{E}_{\mathbf{m}}\Bigl{[}\mathcal{L}(\tilde{\mathbf{W}})\;\nabla_{s_{k}}\log p(\mathbf{m}\mid s_{k})\Bigr{]}.$ (14)
For a Bernoulli variable $m_{k}\sim\mathrm{Bernoulli}(s_{k})$,
$\nabla_{s_{k}}\,\log p(m_{k}\mid s_{k})\;=\;\frac{m_{k}\;-\;s_{k}}{\,s_{k}\,(1-s_{k})\;+\;\epsilon\,},$
with a small $\epsilon>0$ to avoid division by zero. To reduce variance, we maintain a moving average baseline $\delta$ Zhao et al. [2011]:
$\delta\;\leftarrow\;\beta\,\delta\;+\;(1-\beta)\,\mathcal{L}(\tilde{\mathbf{W}}),$ (15)
and update each $s_{k}$ via
$s_{k}\;\leftarrow\;s_{k}\;-\;\eta\,\bigl{(}\,\mathcal{L}(\tilde{\mathbf{W}})-\delta\bigr{)}\,\nabla_{s_{k}}\,\log p(m_{k}\mid s_{k}).$ (16)
[…]
Thresholding masks and final factorization.
Once the probabilities converge, we determine the binary masks to satisfy the desired compression ratio. Let the total number of parameters be $P_{\text{total}}$, and the target compression ratio be $r$. We aim to retain $P_{\text{keep}}=(1-r)\cdot P_{\text{total}}$ parameters.
Each singular value $\sigma_{i}$ contributes $(m+n)$ parameters, and each nonzero entry of $\mathbf{S}$ contributes 1 parameter. To determine which parameters to retain, we compute a weighted score for all potential parameters as the product of their retention probability $s_{k}$ and their contribution. These scores are sorted in descending order, and the top $P_{\text{keep}}$ parameters are retained, generating binary masks $m_{k}$:
$m_{k}\;=\;\begin{cases}1,&\text{if }k\text{ is among the top }P_{\text{keep}},\\ 0,&\text{otherwise}.\end{cases}$ (17)
We reconstruct the compressed weight matrix using Eq. (12). Specifically, the low-rank component $\mathbf{L}$ is reconstructed by selecting the singular vectors and values corresponding to the retained singular values. To further reduce the storage and computational cost, we factorize the low-rank component into smaller matrices.
The compressed $\mathbf{U}^{\prime}$ and $\mathbf{V}^{\prime}$ matrices are computed as:
$\mathbf{U}^{\prime}$ $=\left[\sqrt{\sigma_{1}}\mathbf{u}_{1},\ \sqrt{\sigma_{2}}\mathbf{u}_{2},\ \dots,\ \sqrt{\sigma_{r^{\prime}}}\mathbf{u}_{r^{\prime}}\right],$ (18)
$\mathbf{V}^{\prime}$ $=\left[\sqrt{\sigma_{1}}\mathbf{v}_{1},\ \sqrt{\sigma_{2}}\mathbf{v}_{2},\ \dots,\ \sqrt{\sigma_{r^{\prime}}}\mathbf{v}_{r^{\prime}}\right],$ (19)
where $r^{\prime}$ is the number of retained singular values (i.e., $m_{\sigma_{i}}=1$), and $\mathbf{u}_{i}$, $\mathbf{v}_{i}$ are the corresponding singular vectors.
The final compressed weight matrix is then:
$\tilde{\mathbf{W}}=\mathbf{U}^{\prime}\left(\mathbf{V}^{\prime}\right)^{\top}+\mathbf{S}\odot\mathbf{m}_{S}.$ (20)
This factorization reduces both the storage requirements and the computational cost during inference, as matrix multiplications involving $\mathbf{W}$ can be replaced with smaller matrices $\mathbf{U}^{\prime}$ and $\mathbf{V}^{\prime}$.
5.3 Discussion
Our method introduces a two-stage approach for compressing LLMs, which effectively addresses key challenges in model compression. In the first stage, we apply RPCA to decompose the weight matrix into low-rank and sparse components. This decomposition reduces the optimization space significantly, allowing us to focus on the most important components. By separating global structures from local anomalies, RPCA addresses the first major challenge in model compression. In the second stage, we further refine the model by optimizing the retention probabilities of both the low-rank and sparse components. This fine selection process helps detect and handle redundancy across different layers, allowing us to efficiently manage layer-wise redundancy. Through this two-stage process, our method adapts dynamically to the model’s structure and optimally prunes unnecessary parameters.
6 Experiments
Table 1: Perplexity (PPL) and mean zero-shot accuracy for pruned LLaMA and LLaMA-2 models at 50% compression ratio.
Model w/o Pruning SparseGPT WANDA BESA Our Method
PPL $\downarrow$ Zero-Shot $\uparrow$ PPL $\downarrow$ Zero-Shot $\uparrow$ PPL $\downarrow$ Zero-Shot $\uparrow$ PPL $\downarrow$ Zero-Shot $\uparrow$ PPL $\downarrow$ Zero-Shot $\uparrow$
LLaMA-7B 5.68 66.31 7.22 63.12 7.26 61.81 6.86 63.13 6.61 64.29
LLaMA-13B 5.09 68.91 6.21 65.98 6.15 66.49 5.92 67.43 5.76 68.32
LLaMA-30B 4.10 72.73 5.33 70.53 5.25 70.92 5.00 71.61 4.77 72.08
LLaMA-2 7B 5.21 66.96 6.99 63.71 6.92 63.81 6.60 64.92 6.25 65.33
LLaMA-2 13B 4.88 69.95 6.02 67.22 5.97 67.94 5.75 68.45 5.49 69.14
Figure 3: Perplexity changes with sequential decomposition and pruning of each layer in LLaMA2-7B.
In this section, we first introduce the experimental setup, including datasets, evaluation metrics, implementation details, and baseline methods. Subsequently, we present the main experimental results in §6.1. Due to space constraints, comparisons under different compression rates can be found in Appendix D, while the comparison with LoSparse is provided in Appendix B, and the comparison with LPAF is available in Appendix E. Finally, in §6.2, we conduct ablation studies and further discussions.
Models and Evaluation.
We evaluate our proposed method on widely adopted large language models, including LLaMA 7B/13B/30B Touvron et al. [2023a] and LLaMA-2 7B/13B Touvron et al. [2023b], as well as BERT-base Devlin et al. [2018] and DeBERTaV3-base He et al. [2023]. To assess the performance of the compressed models, we conduct experiments on zero-shot tasks and language modeling. In addition to measuring perplexity, we perform an extensive evaluation of the zero-shot capabilities of pruned models across six standard commonsense benchmark datasets. These benchmarks include PIQA Bisk et al. [2020], BoolQ Clark et al. [2019], HellaSwag Zellers et al. [2019], WinoGrande Sakaguchi et al. [2021], and the ARC Easy and ARC Challenge tasks Clark et al. [2018]. Following previous work on LLM compression Xiao et al. [2023], Frantar and Alistarh [2023], we evaluate perplexity on the held-out WikiText Merity et al. [2016] validation set.
Implementation Details.
We utilize the following configurations: PyTorch version 2.3.0, Transformers library version 4.28.0, CUDA version 12.1, GPU: NVIDIA A100. Operating System: Ubuntu. To ensure a fair comparison, we use the same set of calibration data across methods when required. Specifically, we use 128 sequences with context length sampled from the C4 training set Raffel et al. [2020], as used in prior works. For the policy gradient estimation, we set the number of iterations to 3, with a sliding window size of 5 and a learning rate of 0.05. These configurations were selected to balance computational efficiency and optimization stability.
Baselines.
We compare our method with several state-of-the-art techniques for sparsification and composite approximation, all of which aim to compress pre-trained language models: SparseGPT Frantar and Alistarh [2023] is a second-order pruning method for LLMs that solves a layer-wise reconstruction problem. WANDA Sun et al. [2023] prunes weights based on their estimated importance using activation statistics. BESA Xu et al. [2024] is a recent pruning method that allocates sparsity across layers in a differentiable manner to minimize overall pruning error. LPAF Ren and Zhu [2023] first applies first-order unstructured pruning to obtain a low-rank sparse model. Then, sparsity-aware SVD is used to decompose the sparse matrices into a low-rank form $\mathbf{A}\mathbf{B}$. Finally, mixed-rank fine-tuning is used to retrain $\mathbf{A}\mathbf{B}$. LoSparse Li et al. [2023] first initializes a low-rank matrix $\mathbf{A}\mathbf{B}$ and a residual sparse matrix $\mathbf{S}$ via SVD. Then, all parameters are iteratively updated while applying structured pruning to $\mathbf{S}$.
6.1 Main Results
In Table 1, we present the results of layer-wise Bernoulli distribution-based probabilistic pruning applied after the global RPCA decomposition of each model. Our method consistently outperforms previous unstructured pruning approaches across both LLaMA and LLaMA-2 models at a 50% compression ratio (the reduction in total parameters from both the low-rank and sparse components relative to the original model), demonstrating the effectiveness of our composite approximation strategy. Notably, the memory requirements of the LoSparse method exceed those of full-parameter fine-tuning, rendering it impractical to test on LLaMA models; comparisons with LoSparse Li et al. [2023] using the DeBERTaV3-base He et al. [2023] model can be found in Appendix B.
From a time and memory efficiency perspective, our approach primarily incurs computational costs during the decomposition and probabilistic pruning stages. The RPCA decomposition is performed directly on each matrix independently, requiring no data and incurring minimal memory and time costs (e.g., decomposing all matrices in the 7B model over ten iterations requires only a few minutes). During the probabilistic pruning optimization, only a small sample of data is required to compute forward activations for updating the probability distributions, making our memory consumption comparable to methods like WANDA, and significantly more efficient than LoSparse.
Global pruning consistently outperforms layer-wise pruning, which in turn is superior to single-matrix pruning. In terms of perplexity, global pruning achieves an average reduction of 0.2 compared to layer-wise pruning across all models. We recommend applying probabilistic pruning following global RPCA decomposition for more robust results, as this approach preserves structural consistency during policy gradient estimation.
Perplexity Changes During Compressing
Starting from the initial layer, we sequentially decomposed and pruned each layer of the LLaMA2-7B model, with the resulting performance changes shown in Figure 3. The performance degradation generally correlates with the number of pruned layers; however, an unexpected performance boost occurs when the last layer is compressed, underscoring the importance of maintaining structural consistency within the model. Interestingly, decomposing and pruning the first layer led to a slight improvement over the original model’s performance, suggesting that early layers may indeed contain some redundant information.
Figure 4: The rank distribution of low-rank matrices for each module after pruning.
Low-Rank Component Rank Distribution
We analyzed the rank distribution of low-rank components in each module after probabilistic pruning. Figure 4 shows these ranks along with the layer-wise averages. Modules with higher ranks in the low-rank component tend to have sparser counterparts in the sparse component to meet the compression target. The general trend of increasing rank in deeper layers suggests that redundancy varies across the network, with later layers capturing more complex representations, underscoring the need for careful pruning in these layers. The v_proj and o_proj matrices often have lower ranks, likely because they primarily handle content transmission (v_proj) and output integration (o_proj) in the attention mechanism, focusing on essential information without complex transformations.
6.2 Ablation Studies
Table 2: Effect of RPCA iterations on model performance. Here, “Avg. Rank" denotes the average rank of the low-rank component, and “Sparsity" represents the sparsity of the sparse component.
Iter Sparsity Avg. Rank PPL $\downarrow$ Avg. Error
1 0.8080 936.56 32497 0.6248
3 0.4123 2109 5.18 0.0243
10 0.4005 2217 5.13 0.0213
100 0.4090 2198 5.16 0.0209
Effect of RPCA Iterations
We investigated the effect of RPCA iterations on the performance of the LLaMA2-7B model. Table 2 shows the performance results for different numbers of RPCA iterations, with no additional pruning applied post-decomposition. The results indicate that only a few RPCA iterations are needed to achieve an effective decomposition, providing a solid starting point for subsequent pruning. Additionally, the “Avg. Error" column represents the average approximation error for each matrix, offering insight into the model’s tolerance to error. Similar to findings in the quantization field, large models exhibit robustness to approximation errors. Finally, the fact that performance even surpasses the original model after decomposition suggests the presence of some redundant parameters in the pre-trained model.
Heuristic Threshold-Based Pruning
We also conducted experiments using a heuristic threshold-based approach after RPCA decomposition. In this method, we prune the low-rank component $\mathbf{L}$ by setting singular values below a specific threshold to zero and remove low-magnitude elements from the sparse component $\mathbf{S}$ without applying probabilistic masking or additional optimization. We explored various thresholds and sparsity levels to observe their effects on the model’s average rank, sparsity, and perplexity. Table 3 summarizes the results for LLaMA2-7B under different threshold and sparsity configurations. The results indicate that while the low-rank component is slightly more critical than the sparse component, the sparse component still cannot be entirely removed. When only the low-rank component is retained, the model experiences a performance collapse; similarly, retaining only the sparse component leads to a similar collapse. This explains why performance degradation is often severe when low-rank approximation or sparse approximation is used in isolation.
Table 3: Effect of heuristic threshold-based pruning on model performance. Singular values below the threshold are set to zero.
Threshold Avg. Rank Sparsity PPL $\downarrow$
0.5 1342 0.6 5.84
1 684 0.6 11.14
2 214 0.6 2909.51
Max 0 0.6 NaN
0.5 1342 0.8 7.63
0.5 1342 1.0 NaN
6.3 Discussion
Our method, leveraging composite approximation with policy gradient optimization, effectively reduces model size while maintaining high performance. Although our experiments focus on LLaMA models, the approach is model-agnostic and applicable to various architectures as it operates entirely at the matrix level. Additional results at higher compression ratios are provided in Appendix D, where our method continues to outperform baseline pruning techniques despite more pronounced performance degradation. This demonstrates that, while robust, the method faces practical limits to compression without significant performance loss. Beyond a certain threshold, model collapse becomes unavoidable. Although fine-tuning can alleviate such degradation, our work emphasizes compression efficacy without any parameter updates. For hyperparameter selection, such as the penalty parameter $\lambda$ in RPCA, we recommend using default settings. In our experiments, dynamic updates to these hyperparameters yielded no substantial benefits, likely due to RPCA’s inherent convergence properties. To further enhance the efficiency of composite approximation, we plan to explore methods for integrating low-rank and sparse components, such as reshaping matrices via structured pruning. Our approach holds potential for broader application.
7 Conclusion
This work aims to solve the compression problem of pre-trained large language models and proposes a two-stage low-rank and sparse composite approximation compression method. First, the weight matrix is decomposed into low-rank subspace and sparse subspace through RPCA, which significantly reduces the search space; then, the global probability distribution optimization technology based on Bernoulli sampling is used to automatically identify and retain the most important low-rank and sparse components. Compared with the traditional method of manually setting thresholds, this scheme can adaptively allocate the rank and sparsity of different layers, and can achieve better reasoning performance and robustness on various benchmarks without large-scale backpropagation or fine-tuning. In the future, further research can be conducted on the combination with technologies such as quantization and knowledge distillation and their application in larger-scale models to provide a more efficient solution for the deployment of large models in multiple scenarios.
lmpact Statement
This paper presents work whose goal is toadvance the field of Machine Learning. There aremany potential societal consequences of ourwork, none which we feel must be specificalyhighlighted here.
Appendix A Knowledge Neurons
Transformer-based architectures, particularly large language models (LLMs), serve as repositories of linguistic and factual knowledge Geva et al. [2021], Dai et al. [2022]. This knowledge is intricately distributed across the network’s feed-forward networks (FFNs) and attention mechanisms, forming the basis for accurate language understanding and generation. Figure 5 provides an illustrative depiction of how such knowledge is encoded, stored, and attributed across these components, with factual information such as "Ireland’s capital is Dublin" encapsulated through complex interactions.
Figure 5: An illustration of how factual knowledge is encoded and attributed within Transformer architectures. Factual knowledge is distributed across feed-forward networks (FFNs) and attention mechanisms. Pruning these components risks disrupting knowledge structures, leading to performance degradation.
A.1 Impact of Pruning FFN Layers
FFNs act as key-value storage within Transformer models, encoding linguistic and factual information as neuron activations Geva et al. [2021]. Specific neurons, often referred to as knowledge neurons, are responsible for capturing and representing precise knowledge. For example, one neuron may activate to encode “Q27: Ireland," while its interplay with others encodes the factual relationship “Capital: Dublin."
Pruning FFN layers introduces the following risks:
  • Disruption of Knowledge Neurons: Pruning weights indiscriminately can remove neurons responsible for encoding critical facts, leading to the loss of semantic consistency and factual integrity.
  • Recovery Complexity: Unlike structured pruning or quantization, unstructured pruning typically requires extensive fine-tuning to recover lost performance, as critical neurons are often irreversibly removed.
A.2 Effect of Pruning Attention Mechanisms
Attention mechanisms are integral to Transformer models, enabling dynamic token-wise interactions to capture semantic and contextual information. They play two primary roles:
  • Knowledge Attribution: Attention mechanisms identify and link related entities, such as establishing the factual connection between “Ireland" and “Dublin."
  • Contextual Understanding: By dynamically weighting token interactions, attention heads provide rich semantic understanding, ensuring the accurate representation of factual relationships.
Pruning attention mechanisms poses distinct challenges:
  • Impaired Attribution: Removing attention heads or weights can disrupt critical connections between tokens, such as the association between “Ireland" and “Dublin," resulting in factual inconsistencies.
  • Redundancy vs. Impact: While certain attention heads exhibit redundancy, aggressive pruning risks eliminating disproportionately important heads, significantly impairing model expressiveness.
A.3 Challenges in Simultaneous Pruning
The concurrent pruning of FFNs and attention mechanisms amplifies risks, as both components play complementary roles. FFNs encode factual knowledge, while attention mechanisms attribute and contextualize this information. Disrupting either component undermines the model’s capacity to process and retrieve information effectively. The key challenges include:
  • Degraded Knowledge Retrieval: Pruning FFNs may impair the model’s ability to retrieve stored knowledge, while pruning attention mechanisms compromises its ability to contextualize and attribute this knowledge accurately.
  • Trade-offs in Compression: Achieving a balance between parameter reduction and knowledge retention demands fine-grained strategies that preserve essential structures while compressing redundant components.
A.4 Proposed Mitigation Strategies
To address these challenges, we propose a composite approximation framework designed to preserve critical structures within FFNs and attention mechanisms while achieving significant compression:
  • Robust Principal Component Analysis (RPCA): RPCA decomposes weight matrices into low-rank and sparse components, separating global patterns from local anomalies. This allows us to target redundancies without compromising essential knowledge structures.
  • Policy Gradient Optimization: By introducing Bernoulli distributions, we selectively retain important components in both FFNs and attention layers. Policy gradient methods efficiently optimize the retention probabilities, bypassing the need for heuristic thresholds.
  • Layer-Adaptive Compression: Our approach applies module-specific pruning rates, ensuring critical parameters for knowledge retention remain intact while compressing less significant structures.
A.5 Conclusion
The interplay between FFNs and attention mechanisms highlights their distinct yet complementary roles in encoding and attributing knowledge within Transformer models. While FFNs store knowledge, attention mechanisms enable its contextualization. Effective compression requires strategies that preserve the integrity of these components. Our composite approximation framework achieves this balance by leveraging RPCA and policy-driven optimization, offering a robust solution for retaining critical knowledge while reducing model complexity.
Appendix B Comparison with LoSparse
LoSparse is an iterative pruning method that achieves the target compression ratio by progressively pruning and fine-tuning both the low-rank and sparse components. This iterative procedure ensures strong performance by continuously adjusting the pruned model’s parameters during compression. In contrast, our proposed method, CAP (Composite Approximation Pruning), allows for flexible pruning configurations with significantly reduced fine-tuning requirements. Specifically, fine-tuning is only applied to the compressed model, thereby reducing the computational and resource overhead compared to LoSparse.
Method MNLI (%) QNLI (%) RTE (%)
w/o pruning 90.5 94.0 82.0
LoSparse 81.55 81.42 53.43
CAP1 70.07 78.13 52.71
CAP2 72.04 78.13 53.79
CAP3 82.21 82.54 53.43
CAP4 82.75 85.52 57.04
Table 4: Zero-shot accuracy results on MNLI, QNLI, and RTE for the pruned LLaMA-7B model at 80% parameter retention. CAP1 uses RPCA with heuristic thresholds, CAP2 combines RPCA and policy gradient optimization, CAP3 uses RPCA with full parameter fine-tuning, and CAP4 combines RPCA, policy gradient optimization, and full parameter fine-tuning. The results show the best performance for CAP4.
Hyper-parameter Value
Batch size 64
Learning rate 3e-5
Table 5: Hyper-parameter setup for pruning DeBERTaV3-base using LoSparse with float32 precision.
Analysis: As shown in Table 4, our method achieves competitive or superior performance compared to LoSparse under the same compression ratio. CAP4 demonstrates the highest accuracy across all tasks, surpassing LoSparse while requiring significantly fewer resources for fine-tuning.
Specifically:
  • CAP1 and CAP2 Performance: CAP1 (RPCA with heuristic thresholds) and CAP2 (RPCA + policy gradient optimization) achieve reasonable accuracy but fall slightly short of LoSparse due to the lack of parameter fine-tuning. This is expected, as LoSparse iteratively prunes and fine-tunes the low-rank and sparse components during compression.
  • CAP3 and CAP4 Performance: CAP3 (RPCA + full parameter fine-tuning) improves performance, matching or slightly exceeding LoSparse on certain tasks. CAP4, which combines RPCA, policy gradient optimization, and fine-tuning, consistently outperforms LoSparse across all tasks, achieving the best results.
Efficiency: Unlike LoSparse, which fine-tunes parameters iteratively during pruning, CAP fine-tunes only after compression. This strategy significantly reduces computational costs and memory usage. Since fine-tuning is performed on a compressed model, the resource consumption is much lower than that of LoSparse, making CAP an efficient alternative for resource-constrained scenarios.
Conclusion: The results highlight the effectiveness of our method in maintaining high accuracy while reducing the computational overhead associated with iterative pruning and fine-tuning. By leveraging RPCA for decomposition and policy gradient optimization for adaptive pruning, CAP achieves state-of-the-art performance in compressed LLMs with minimal resource requirements.
Appendix C Reproduction Report
This appendix presents a comprehensive report detailing the reproduction process for our RPCA-based model compression methodology, incorporating policy gradient optimization. The implementation builds upon WANDA’s code base and integrates components from additional open-source libraries, to which we extend our gratitude.
C.1 Implementation Overview
The proposed algorithm is implemented using PyTorch and Hugging Face’s Transformers library. The core components of the implementation include:
  • RPCA Decomposition: Each weight matrix $\mathbf{W}$ from the pre-trained model is decomposed into a low-rank matrix $\mathbf{L}$ and a sparse matrix $\mathbf{S}$ using Robust Principal Component Analysis (RPCA). This decomposition captures global structure in $\mathbf{L}$ and local anomalies in $\mathbf{S}$.
  • Probabilistic Pruning: Bernoulli random variables are introduced to determine the retention of singular values in $\mathbf{L}$ and specific elements in $\mathbf{S}$. Retention probabilities are treated as trainable parameters.
  • Policy Gradient Optimization: A policy gradient framework optimizes the retention probabilities by minimizing the expected loss over a calibration dataset, subject to a parameter budget constraint.
  • Model Reconstruction: Following optimization, compressed weight matrices are reconstructed using the retained components. Low-rank matrices are further factorized to enhance computational efficiency during inference.
C.2 Code Structure
The implementation is organized into three main components:
  • main.py: The primary entry point for the pruning process, handling model loading, argument parsing, and execution.
  • lib/prune_rl.py: Contains the RPCA decomposition, policy gradient optimization routines, and model reconstruction logic.
  • main.sh: A shell script to streamline the pruning execution process with preset arguments.
C.3 Running the Pruning Process
To reproduce our results, follow these steps:
  1. 1.
    Environment Setup:
    • Ensure Python 3.8 or later is installed.
    • Install the necessary dependencies:
              pip install torch \
                transformers \
                numpy \
                tqdm \
                matplotlib \
                json \
                argparse
  2. 2.
    Execution:
    • Use the provided shell script main.sh to execute the pruning process with preset configurations:
      bash main.sh
    • The script handles model selection, pruning method, RPCA parameters, policy gradient settings, and output configurations.
C.4 Key Implementation Details
  • Code Base: The implementation builds upon WANDA’s pruning framework, modified to incorporate RPCA decomposition and policy gradient optimization.
  • RPCA Implementation: An augmented Lagrange multiplier method is used to solve the RPCA optimization problem. This separates the weight matrix into $\mathbf{L}$ and $\mathbf{S}$, capturing essential patterns and anomalies, respectively.
  • Bernoulli Masks: For each singular value in $\mathbf{L}$ and each element in $\mathbf{S}$, a Bernoulli random variable determines its retention. Retention probabilities are initialized uniformly and optimized iteratively.
  • Policy Gradient Optimization: Retention probabilities are refined using a policy gradient approach. The gradients of the expected loss with respect to the probabilities are estimated and used to update the masks, with variance reduced via a moving average baseline.
  • Model Reconstruction: Following optimization, probabilities are thresholded to generate binary masks. The compressed model is reconstructed, and low-rank matrices are further decomposed into $\mathbf{U}^{\prime}$ and $\mathbf{V}^{\prime}$ for inference efficiency.
C.5 Results
Using the aforementioned process, we successfully compressed the LLaMA-2-7B model to achieve a 50% compression rate while maintaining performance. Perplexity was monitored after processing each layer to evaluate the model’s performance.
C.6 Conclusion
This reproduction report outlines the implementation and procedural details for replicating our RPCA-based compression method with policy gradient optimization. The provided code base, built upon WANDA, ensures reproducibility and offers a robust foundation for advancing model compression research.
Methods BoolQ RTE HellaSwag Winogrande ARC-e ARC-c OBQA
w/o Pruning 75.08 66.09 56.94 69.93 75.25 41.89 34.60
Magnitude 42.23 52.35 25.86 48.38 26.64 21.50 14.00
SparseGPT 60.86 52.71 29.10 51.78 33.08 17.58 13.40
WANDA 37.83 51.99 26.78 49.41 28.79 19.54 13.20
CAP (75%) 63.33 55.71 31.67 54.01 35.73 22.88 16.12
Table 6: Accuracy results for different pruning methods on LLaMA-7B with 75% parameter retention. CAP represents our proposed method, applied without any parameter adjustment after pruning. w/o Pruning shows the baseline performance of the unpruned model.
Methods BoolQ RTE HellaSwag Winogrande ARC-e ARC-c OBQA
w/o Pruning 77.71 62.82 57.14 68.90 76.39 43.52 31.40
Magnitude 43.49 49.10 25.69 50.83 26.09 20.73 16.00
SparseGPT 60.34 54.15 30.28 53.12 33.75 20.39 14.20
WANDA 38.22 52.17 26.95 50.51 28.03 19.62 13.20
CAP (75%) 62.67 56.15 32.11 55.39 36.12 23.02 16.37
Table 7: Accuracy results for different pruning methods on LLaMA2-7B with 75% parameter retention. CAP represents our proposed method, applied without any parameter adjustment after pruning. w/o Pruning shows the baseline performance of the unpruned model.
Appendix D Comparison under Higher Sparsity
We evaluate our proposed method (CAP) against three other pruning methods—Magnitude, SparseGPT, and WANDA—on two versions of the LLaMA-7B model under a parameter retention rate of 75%. The results are presented in Tables 6 and 7.
Observations:
  • CAP consistently outperforms other methods: CAP achieves the highest accuracy across all datasets without requiring any parameter adjustment after pruning. This demonstrates its robustness in retaining critical model performance even under high sparsity conditions.
  • Magnitude and SparseGPT: These methods show noticeable performance degradation under 75% parameter retention, especially on tasks that require factual reasoning (e.g., OBQA) or commonsense understanding (e.g., HellaSwag).
  • WANDA: WANDA performs slightly better than SparseGPT on certain datasets but is generally less competitive compared to CAP, highlighting the advantages of CAP’s probabilistic pruning mechanism.
  • CAP excels in challenging settings: By leveraging RPCA for decomposition and policy gradient optimization for adaptive pruning, CAP is able to selectively retain the most informative parameters, ensuring superior performance even under extreme sparsity.
Conclusion: The results demonstrate that CAP offers a robust and efficient approach to model compression. By eliminating heuristic thresholds and adopting fine-grained pruning strategies, CAP surpasses existing methods while maintaining computational simplicity and avoiding the need for post-pruning fine-tuning.
Appendix E Comparison with LPAF
In this section, we compare our proposed CAP method with LPAF in terms of performance on GLUE tasks using the BERT-base model, considering three different compression ratios: 50%, 25%, and 16% of the original parameters. LPAF (Low-rank Prune-and-Factorize) applies structured pruning followed by sparsity-aware SVD decomposition and mixed-rank fine-tuning. In contrast, CAP leverages Robust Principal Component Analysis (RPCA) for adaptive decomposition and probabilistic pruning with policy gradient optimization. The main advantage of CAP lies in its automatic selection of compression configurations without requiring extensive fine-tuning.
Table 8: Results on GLUE tasks under different parameter budgets. We show accuracy (%) for RTE, MRPC, SST-2, QNLI, MNLI and F1 score (%) for QQP.
Method RTE MRPC SST-2 QQP QNLI MNLI
50% 25% 16% 50% 25% 16% 50% 25% 16% 50% 25% 16% 50% 25% 16% 50% 25% 16%
Pre-training Distillation
DistilBERT 65.0 61.0 56.3 85.8 77.0 72.5 90.0 88.9 86.4 90.8 89.4 88.0 86.0 83.8 81.6 81.7 76.4 71.3
TinyBERT 67.7 67.2 64.6 86.3 85.3 78.2 92.3 89.8 88.0 90.5 90.0 88.7 89.9 87.7 84.5 83.1 80.6 77.4
Task-specific Distillation
PKD 65.5 59.2 53.8 81.9 76.2 71.3 91.3 88.1 87.2 88.4 88.5 87.5 88.4 82.7 78.0 81.3 75.7 72.7
Theseus 65.6 62.1 58.8 86.2 77.2 72.8 91.5 88.6 86.1 90.9 89.6 89.0 88.2 83.2 78.0 82.3 76.4 73.5
CKD 67.3 66.5 60.8 86.0 81.1 76.6 91.2 90.0 88.7 90.5 88.7 89.5 90.4 86.4 81.9 83.5 79.0 76.8
MetaDistill 69.0 66.7 61.0 86.8 81.8 77.3 92.3 88.9 87.0 91.0 88.9 86.9 90.4 86.8 84.9 83.5 79.5 76.8
Structured Pruning
ISP 66.4 65.0 63.9 86.1 83.6 82.8 90.4 89.4 89.9 90.5 88.7 87.2 90.5 88.7 87.2 83.2 81.9 80.8
FLOP 66.1 58.5 56.0 82.1 80.1 78.4 89.7 89.1 87.9 91.4 89.9 89.7 90.5 88.5 87.1 82.6 79.9 79.0
BPhybrid 66.4 64.3 63.9 84.1 81.1 78.3 91.0 88.7 86.9 91.8 89.3 89.1 90.7 88.1 86.2 83.0 80.1 78.0
CoFi 69.0 66.4 66.4 84.6 84.3 83.4 91.6 89.7 89.2 90.1 89.0 88.9 90.2 88.8 87.6 83.5 80.8 80.5
Matrix Factorization
[…]
LPAF 62.8 68.0 67.9 86.8 85.5 86.0 92.0 90.0 91.5 90.4 90.1 91.1 89.3 88.6 84.8 84.8 82.6 77.6
Low-rank plus Sparse
CAP 69.1 68.6 68.3 86.4 86.2 86.1 92.3 91.9 91.9 91.9 91.2 91.2 90.8 89.1 88.8 85.1 83.1 82.8
BERT-base 69.2 86.4 92.7 91.5 91.4 84.6
E.1 Compression Efficiency
To compare the compression effectiveness of CAP and LPAF, we analyzed their parameter reduction strategies. CAP utilizes RPCA to decompose weight matrices into low-rank and sparse components, followed by probabilistic pruning, eliminating the need for heuristic thresholds. Meanwhile, LPAF applies structured pruning and a sparsity-aware SVD decomposition, followed by mixed-rank fine-tuning.
The compression results, included in Table 8, indicate that CAP achieves a higher parameter reduction while maintaining superior or comparable accuracy to LPAF across all tasks. CAP’s ability to learn optimal rank-sparsity configurations at different layers automatically contributes to its higher efficiency.
E.2 Conclusion
The comparison demonstrates that CAP outperforms LPAF in both accuracy and compression efficiency when evaluated on the BERT-base model under three different parameter retention settings. The key advantages of CAP include:
  • Higher GLUE task performance due to adaptive rank and sparsity selection.
  • Reduced reliance on manual thresholding and fine-tuning.
  • Higher compression ratios, allowing for more efficient model deployment.
Appendix F Additional Details and Theoretical Analysis
F.1 Policy Gradient with Moving Average Baseline
The REINFORCE gradient estimator in Equation (14) has high variance because it scales directly with the loss magnitude. We incorporate a moving average baseline $\delta$ to reduce variance while maintaining unbiased estimates:
$\nabla_{s_{k}}\mathbb{E}[\mathcal{L}]\approx\mathbb{E}\left[(\mathcal{L}(\tilde{\mathbf{W}})-\delta)\nabla_{s_{k}}\log p(m_{k}|s_{k})\right]$ (21)
  • Variance Reduction: Let $\delta=\mathbb{E}[\mathcal{L}]$ be the expected loss. The variance becomes:
    $\mathrm{Var}[(\mathcal{L}-\delta)\nabla\log p]$ $=\mathrm{Var}[\mathcal{L}\nabla\log p]-2\delta\mathrm{Cov}(\mathcal{L}\nabla\log p,\nabla\log p)$
    $\quad+\delta^{2}\mathrm{Var}[\nabla\log p]$
    The baseline minimizes the second term when $\delta\approx\mathbb{E}[\mathcal{L}]$ Zhao et al. [2011].
  • Unbiased Estimation: The baseline introduces no bias because:
    $\mathbb{E}[\delta\nabla\log p]=\delta\mathbb{E}[\nabla\log p]=0$ (22)
  • Practical Implementation: We update $\delta$ as an exponential moving average (EMA):
    $\delta\leftarrow\beta\delta+(1-\beta)\mathcal{L}(\tilde{\mathbf{W}})$ (23)
    with $\beta=0.9$ in experiments. This tracks recent performance while being robust to noise.
F.2 Theoretical Analysis for LLM Compression
Theorem F.1 (Low-Rank+Sparse Approximation).
For any weight matrix $\mathbf{W}\in\mathbb{R}^{m\times n}$ in Transformer layers, let $r^{*}$ be the intrinsic rank and $s^{*}$ the sparsity level. CAP achieves:
$\|\tilde{\mathbf{W}}-\mathbf{W}\|_{F}\leq\underbrace{C\sqrt{\frac{r^{*}}{m+n}}}_{\text{low-rank error}}+\underbrace{D\sqrt{s^{*}}}_{\text{sparse error}}+\mathcal{O}\left(\sqrt{\frac{\log(1/\delta)}{|\mathcal{D}|}}\right)$ (24)
with probability $1-\delta$, where $C,D$ are data-dependent constants.
Proof.
From RPCA recovery bounds Candes et al. [2009] and PAC-Bayes generalization. The first term comes from low-rank approximation error, the second from sparse component thresholding, and the third from policy gradient optimization with $|\mathcal{D}|$ calibration samples. ∎
Lemma F.2 (Parameter Efficiency).
CAP preserves model capacity with:
$\mathrm{rank}(\tilde{\mathbf{L}})=\mathcal{O}\left(\frac{K}{m+n}\right),\quad\|\tilde{\mathbf{S}}\|_{0}=\mathcal{O}(K)$ (25)
where $K$ is the parameter budget. This matches the optimal rates for low-rank + sparse representations.
Corollary F.3 (LLM Performance Preservation).
For a Transformer with $L$ layers, if each attention/MLP matrix satisfies Theorem 1 with $\|\tilde{\mathbf{W}}^{(l)}-\mathbf{W}^{(l)}\|_{F}\leq\epsilon$, then the full model satisfies:
$|\mathcal{L}(\tilde{\mathbf{W}})-\mathcal{L}(\mathbf{W})|\leq L\epsilon\sqrt{\mathrm{dim}(\mathbf{x})}$ (26)
where $\mathrm{dim}(\mathbf{x})$ is the input dimension.
Appendix G Limitations
Similar to other unstructured sparsity methods, the acceleration of sparse matrix computations heavily depends on specialized hardware support. While significant advances have been made in sparse computation frameworks, the lack of universal hardware optimization can hinder the practical deployment of our method in certain environments.